I'm sorry, I forgot to include the total.
A student is analyzing the relationship between gender and favorite type of book out of mystery, sci-fi and romance. Calculate whether gender and type of favorite book are independent at the .05 level of significance. Show all five steps of this test. Here are the random samples:
Mystery--Sci-Fi-- Romance-- Total
Male 28-- 152-- 218-- 398
Femal 213-- 102-- 189-- 504
Total 241-- 254-- 407-- 902

Answer by ewatrrr(24785)

Ho: Gender is not related to a person's favorite type of book
Ha: Gender is related to a person's favorite type of book
|
Level of significance = .05
|
Test Statistic is Chi-square
Χ^2 = Σ [ (Or,c - Er,c)2 / Er,c ] where Er,c = (nr * nc) / n
To find X^2 recommend using calculator found at www.vassarstats.net/newcs.html
X^2 = 143.45
|
DF = 2
(2-1)(3- 1) = 2 (2 rows, 3 columns)
|
p = P(X > 143.45), df= 2
TI calculator
p = X^2cdf ((lowerbound, upperbound, df) = p = X^2cdf ((0,143.45, 2) = .0001
|
.0001 < .05 , reject Ho, Gender is related to a person's favorite type of book